rvc (rvc):

If y(x) is the solution of the differential equation (x+2)dy/dx=x^2+4x-9,x not equal to -2 and y(0),then y(-4) is equal to ?

rvc (rvc):

is it equal to -1?

OpenStudy (perl):

there is a typo in your question, or an omission. what does y(0) = ?

rvc (rvc):

oops! sorry y(0)=0 :)

OpenStudy (perl):

$$ \Large { (x+2) \frac{dy}{dx} = x^2 + 4x - 9 \\ \iff \\ \frac{dy}{dx} = \frac{x^2 + 4x - 9 }{x+2} \\ \iff \\ dy = \frac{x^2 + 4x - 9 }{x+2} ~dx \\ \iff \\ \int dy = \int \frac{x^2 + 4x - 9 }{x+2}~ dx \\ \iff \\ y = \frac 12 x^2+2x-13\ln|x+2|+ C \\ 0 = \frac 12~ 0^2+2(0)-13\ln|0+2|+ C \\ \therefore \\~\\ \\ C = 13 \ln2 \\~\\ \therefore \\ y = \frac 12 x^2+2x-13\ln|x+2|+ 13 \ln 2 \\ \therefore \\ \large {y(-4) = \frac 12 (-4)^2+2(-4)-13\ln|(-4)+2|+ 13 \ln 2 \\= \frac 12 \cdot 16 - 8 - 13 \ln | -2 | + 13\ln 2 \\=8 - 8 - 13 \ln (2) + 13 \ln 2 \\= 0 - 0 = 0 }} $$

rvc (rvc):

so its equal to 0?

OpenStudy (perl):

I think so

OpenStudy (perl):

if we use the absolute values when we integrate. if we don't use the absolute values, then we get an imaginary answer

rvc (rvc):

hhmm.. @rational pls help

OpenStudy (rational):

\[\dfrac{dy}{dx}=\dfrac{x^2+4x-9}{x+2}\] \[y(t)-y(0) = \int\limits_0^t\dfrac{x^2+4x-9}{x+2} dx\]

rvc (rvc):

hmm.. what next

OpenStudy (rational):

are you given the value of \(y(0)\) ?

OpenStudy (rational):

it must be given, double check the question

rvc (rvc):

yes i mentioned it later it's y(0)=0

rvc (rvc):

hmm..

OpenStudy (rational):

then i think y(-4) is undefined as the domain of solution is (-2, infty)

rvc (rvc):

@Michele_Laino pls help

rvc (rvc):

0,2,1,-1 these are the options

OpenStudy (perl):

the domain of the solution (-infinity, -2) U (-2, infinity) the answer is zero, assuming you copied the question correctly

rvc (rvc):

oh okay

OpenStudy (perl):

note that the absolute value takes care of the case when x < - 2, so we don't have a problem. in general : $$ \Large \int \frac 1 x dx = \ln | x | + c $$

OpenStudy (perl):

I can add a step here

OpenStudy (perl):

$$\large \rm y = \begin{cases} \frac 12 x^2+2x-13\ln(x+2)+ 13 \ln 2 ~if ~x > -2\\ \frac 12 x^2+2x-13\ln(-(x+2))+ 13 \ln 2 ~if~ x < -2 \end{cases} $$

OpenStudy (rational):

the inititial value x=0 is on the right side of pole at x=-2 and the value they asked to evaluate is on the left side of pole. So y(-2) does not exist for the solution of this IVP.

OpenStudy (rational):

http://apcentral.collegeboard.com/apc/members/repository/ap07_calculus_DE_domain_fin.pdf explains clearly about domain of solution of IVP

OpenStudy (perl):

@rational "the inititial value x=0 is on the right side of pole at x=-2 and the value they asked to evaluate is on the left side of pole. So y(-2) does not exist for the solution of this IVP." You mean y(-4) does not exist for the solution of this IVP ?

OpenStudy (rational):

Ahh yes y(-4)

OpenStudy (perl):

we could draw a slope field and verify

OpenStudy (rational):

domain of solution for this IVP is x>-2

OpenStudy (perl):

i have to read that pdf, but i dont see a problem with it superficially

OpenStudy (rational):

the attached pdf explains nicely about the domain of solutions and other technical issues

OpenStudy (perl):

i will do that and get back to you

rvc (rvc):

i cant open the pdf

OpenStudy (perl):

@IrishBoy123 any comments

OpenStudy (perl):

domain subtleties

OpenStudy (irishboy123):

original diff eqn = \[\frac{dy}{dx} = \frac{(x+2)^2 - 13}{(x+2)}\] \[= (x + 2) - \frac{13}{x+2}\] only problem is at x = -2, otherwise looks continuous, the || in the ln takes care of when x < -2, hence y(-4) = 0 IMHO

rvc (rvc):

okay thank u

OpenStudy (michele_laino):

I got this solution: \[y\left( x \right) = C + \frac{{{x^2}}}{2} + 2x - 13\ln \left| {x + 2} \right|\]

OpenStudy (rational):

The solution of given IVP is "only" the right part : \[y = \frac 12 x^2+2x-13\ln(x+2)+ 13 \ln 2 ;~~x\gt -2\] you cannot have two different solution curves a IVP contradicting the existence and uniqueness thm.

OpenStudy (rational):

It might seem reasonable to evaluate the integral and simply plugin \(x=-4\). but it wont work here because the domain of IVP is \(x\gt-2\)

OpenStudy (irishboy123):

@rational just to be clear, are you saying we know nothing about y when x < -2, other than its derivative? can we not infer something from the (skewed) symmetry as x+ -> -2 and x- -> -2? just asking, out of my depth right now.

OpenStudy (rational):

yes y(-4) can be anything for other IVP. it is undefined as far as the IVP in the main question is concerned.

OpenStudy (rational):

the pole at x = -2 and the initial value of (0,0) shrink the domain to (-2, infty)

rvc (rvc):

well out of the four options 0 is the ans right :)

rvc (rvc):

@Haseeb96

OpenStudy (haseeb96):

yes

rvc (rvc):

the answer is 0?

OpenStudy (haseeb96):

o par solve kar tu rakha hai ye question .

OpenStudy (haseeb96):

yes, answer is 0